OpenStudy (anonymous):

Solve 3(m-2)-5=8-2(m-4)

OpenStudy (anonymous):

3(m-2)-5=8-2(m-4) 3m -6 -5 = 8 -2m +8 3m + 2m = 8+ 8+11=27 m=27/5

OpenStudy (anonymous):

@eliassaab Omg thank you so much!